# Monte Carlo Physics-informed neural networks for multiscale heat conduction via phonon Boltzmann transport equation

### v3

@@ Abstract
The phonon Boltzmann transport equation (BTE) is widely used for the description of multiscale heat conduction (from $nm$ to $\mu m$ or $mm$) in solid materials. Developing numerical approaches to solve this equation is challenging since it is a 7-dimensional integral-differential equation. In this work, we propose the Monte Carlo physics-informed neural networks (MC-PINNs), which do not suffer from the “curse of dimensionality”, to solve the phonon Boltzmann transport equation for modeling the multiscale heat conduction in solid materials. In MC-PINNs, we utilize a deep neural network to approximate the solution to the BTE and encode the BTE as well as the corresponding boundary/initial conditions using the automatic differentiation. In addition, we propose a novel two-step sampling approach to address the issues of inefficiency and inaccuracy issues in the widely used sampling methods in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial space (Step I) and then draw another number of points randomly in the solid angular space (Step II). The training points are constructed based on the data drawn from the above two steps using the tensor product. The two-step sampling strategy enables the MC-PINNs (1) to model the heat conduction from ballistic to diffusive regimes, and (2) is memory more efficient compared to the conventional numerical solvers or existing PINNs for BTE. A series of numerical examples including a quasi-one-dimensional (quasi-1D) steady/unsteady heat conduction in a film, and the heat conduction in a quasi-two- and three-dimensional square domain, are conducted to justify the effectiveness of the MC-PINN for heat conduction spanning diffusive and ballistic regimes. Finally, we perform a comparison of the computational time and the memory usage of the MC-PINN with one of the state-of-art numerical methods to demonstrate the potential of the MC-PINN for large-scale problems in real-world applications.
@@ 1 Introduction
Microdevices, e.g., microprocessors and microelectronic circuits, often generate heat with great density during operation. Efficient thermal management is crucial to the reliability and optimal performance of these devices. As pointed out in [1], the heat conduction in microdevices like chips is inherently a multiscale problem spanning nanoscale (1-100 nm) to microscale (0.1-100 $\mu$m) or macroscopic (0.1-100 mm). The classical Fourier’s law is a widely used empirical macroscopic model for heat conduction in solids. However, it is not capable of describing the heat conduction in the non-diffusive or ballistic regime [2, 3, 4]. Further, the phonon Boltzmann transport equation (BTE) is generally employed for describing the heat conduction in solid materials spanning the ballistic and diffusive regimes [5, 1, 6, 7].
@@
To advance the understanding of multiscale heat conduction in solid materials, lots of efforts have been put into developing numerical methods for solving the phonon BTE. As mentioned, developing effective numerical methods for efficiently solving the phonon BTE is challenging since it is a 7-dimensional integral differential equation in real-world applications. The conventional numerical methods can be broadly categorized into two main approaches, i.e., the stochastic methods, e.g., direct simulation Monte Carlo (DSMC) [8, 9, 10, 11], etc., and deterministic approaches, e.g., discrete velocity/ordinate method [12, 13, 14], synthetic iterative scheme [15, 16, 17], discrete unified gas kinetic scheme (DUGKS) [18, 19], and so on. The former uses discrete meshes in the temporal-spatial space and is efficient for steady problems in the ballistic regime. However, it converges slowly for problems in the (near) diffusive regime as well as non-stationary problems. The latter employs the discrete meshes in the temporal-spatial-angular space and is able to achieve good accuracy for steady/unsteady problems spanning a wide range of regimes. In general, the number of discrete points in the solid angular space required to obtain accurate results for the heat conduction problem in the diffusive regime is much smaller than in the ballistic regime. The requirement of large numbers of discrete points in the solid angular space for accurately capturing the non-Fourier’s behavior in the ballistic regime leads to intensive computational cost and prohibitive memory usage for large-scale thermal engineering problems in practice.
@@
Deep neural networks (DNNs), which are able to break the curse of dimensionality, have recently emerged as an effective tool for solving partial differential equations (PDEs), such as physics-informed neural networks (PINNs) [20, 21, 22, 23] and DeepRitz [24]. Generally, the DNN-based PDE solvers are mesh-free, which saves significant effort for generating the meshes in conventional methods, e.g., finite volume method, and finite element method. Among the aforementioned deep learning approaches, the PINNs, which encode the PDEs via the automatic differentiation technique, are one of the most widely used methods for solving PDE problems due to their effectiveness and straightforward implementations. Recently, the PINNs have been successfully employed to simulate flows ranging from continuous to rarefied flows by solving the Boltzmann-BGK or Boltzmann equation, demonstrating their capability of handling multiscale problems [25, 26]. Inspired by the aforementioned work for modeling multiscale flows using PINNs, Li et al. proposed to solve the phonon BTE using PINNs to study the heat conduction in solid materials [27, 28]. The results show that the PINNs are capable of modeling heat conduction ranging from ballistic to diffusive regimes with satisfactory accuracy. In particular, the temporal-spatial, as well as the solid angular space, are discretized as a priori for training the PINNs in previous works [25, 27, 29]. For instance, the angular space is discretized utilizing the points in the Gauss Legendre quadrature rule, which is similar to the way widely used in the conventional deterministic approaches for solving phonon BTE, e.g., DUGKS. Although the effort of generating the meshes in temporal-spatial space can be saved for the PINN approaches, it shares the same defect as in the conventional deterministic approaches, i.e., a large number of discrete points in the phase space is required for heat conduction at the ballistic regime. Moreover, for scenarios in real-world applications where the heat conduction at the micro- and nanoscale coexists, a large number of discrete points in the phase space of phonon BTE is also required for accurate descriptions of the non-Fourier’s behavior at the ballistic regime since it is challenging to develop adaptive discretization for heat conduction at different scales. Therefore, we still have difficulty employing the PINNs for heat conduction in large-scale problems due to the expensive computational as well as memory cost. Numerical approaches for solving the phonon BTE that are more flexible as well as efficient are herein still desirable.
@@
In this study, we employ the PINNs as the backbone to solve the phonon BTE due to their effectiveness and capability for handling high-dimensional problems. In particular, we propose a novel two-step sampling approach for the training of PINNs to address the issue of expensive computational and memory costs in the existing PINN models. We refer to the PINN equipped with the proposed two-step sampling approach as the Monte Carlo PINNs, which (1) is capable of modeling the heat conduction spanning ballistic to diffusive regimes, (2) is mesh-free in the temporal-spatial and the angular space, (3) is a unified approach at all regimes without the requirement to employ different discretizations in the angular space for heat conduction at different scales, and (4) requires much less discrete points in the angular space at each training step, and thus is more computationally and memory efficient comparing to the approach in [27].
@@ 2.1 Phonon Boltzmann transport equation
The phonon Boltzmann transport equation (BTE) is a fundamental tool for understanding heat conduction in solid materials at different scales. Various models of the phonon BTE have been developed to address specific physical scenarios [9]. For instance, the full/linearized phonon BTE considers the detailed phonon dispersion and full scattering kernel, enabling the study of complex interactions among different phonon modes [30]. For scenarios that demand frequency-specific considerations, the frequency-resolved BTE under relaxation time approximation is often employed, in which the relaxation rate on the diagonal of phonon scattering matrix is accounted [10, 11, 16, 17]. A widely used model among these approaches is the gray model, where nonlinear phonon dispersion and polarization are not accounted for and the frequency-independent assumption is used [19, 15, 17]. In this study, we focus on the phonon gray model, which is apt for probing multiscale heat conduction in three-dimensional materials subject to minor temperature variations, i.e., $\Delta T\ll T_{ref}$. Based on a certain number of reference variables, e.g., reference temperature $T_{ref}$, specific heat $C_{ref}$, reference length $L_{ref}$, group velocity $v_{g}$ and time $t_{ref}=L_{ref}/v_{g}$, then the dimensionless phonon BTE in the gray model can be obtained as
@@
where $e^{\prime\prime}$ represents the phonon distribution function of energy density, which is a function of time $t$, spatial coordinate $\bm{x}$, and the unit directional vector $\bm{s}=(\cos\theta,\sin\theta\cos\varphi,\sin\theta\sin\varphi)$. Specifically, $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$ are the polar and azimuthal angles, respectively. In addition, $e^{eq}$ denotes the equilibrium distribution function for the energy density; $C_{V}$ and $T$ are the specific heat and temperature, respectively. Moreover, $\text{Kn}=\lambda/L_{ref}$ is the Knudsen number defined as the ratio between the phonon mean free path to the characteristic length for a specific system. Due to the law of energy conservation, we can obtain the following equation for the right-hand side of Equation (1a):
@@
where $d\Omega$ represents the integral over the entire solid angular space. The quantities of interest (QoI) in heat conduction problems including temperature $T$ and heat flux $\bm{q}$ can be obtained by taking the moment of phonon density distribution function, i.e.,
@@
We now introduce the boundary conditions briefly for solving Eq. (1), including the thermalization boundary condition, the periodic boundary condition, diffuse reflection, and so on. The periodic boundary condition, applied in repeated structures, ensures the continuity of the phonon flux and temperature field across opposite boundaries. Diffuse reflection conditions model rough surfaces, where phonons scatter in all directions based on the incoming distribution. In this work, we adopt the thermalization boundary condition for heat conduction problems [31], which allows us to simulate the heat conduction processes accurately for the boundary with a fixed temperature. Under this condition, phonons are absorbed upon striking the boundary, and new phonons are emitted with equilibrium states determined by the boundary temperature $T_{w}$. This is equivalent to implementing Dirichlet boundary conditions for the phonon energy density $e^{\prime\prime}$ at the boundary, with the rest derived from the governing equation:
$e^{\prime\prime}(t,x_{0},\bm{s})=\frac{C_{V}T_{w}}{4\pi},\quad\bm{s}\cdot\bm{n}>0,$ (4)
where $\bm{n}$ is the unit normal vector pointing from the boundary to the computational domain. For the initial conditions, we assume that the distribution function is under equilibrium state at the initial temperature at $t=0$:
$e^{\prime\prime}(t_{0},x,\bm{s})=\frac{C_{V}T_{0}}{4\pi}.$ (5)
Figure 1: Schematic of flow regimes based on the Knudsen number. $L$ is the characteristic length of the system and $\lambda$ is the phonon mean free path.
We would like to discuss that the multiscale heat conduction is generally divided into different regimes based on the Knudsen number, seen in Figure 1. When $\rm{Kn}<0.01$, the characteristic length of the system is much larger than the phonon mean free path, and phonon transport is in the diffusive regime. The resulting temperature field is linear, consistent with the predictions of Fourier’s law of heat conduction. As $\rm{Kn}$ increases, Fourier’s law begins to fail. When $\rm{Kn}>10$, heat transfer is dominated by ballistic effects.
@@ 2.2.1 Physics-informed neural networks
The PINNs for solving phonon BTE are illustrated in Fig. 2. As shown, we first utilize deep neural networks (DNNs), which take inputs as $(t,\bm{x},\bm{s})$, to approximate the solution to Eq. (1), i.e., $e^{\prime\prime}$, and then the governing equation, as well as the boundary and/or the initial conditions, are encoded in the DNNs via the automatic differentiation [32]. The loss function for training the PINNs is composed of the mean squared errors for the residual of the equation, mismatches between the predictions of DNNs and the boundary/initial conditions, which is expressed as follows:
$L=L_{Eq}+L_{BC}+L_{IC},$ (6)
where $L$ is the total loss, $L_{Eq}$, $L_{BC}$, and $L_{IC}$ denote the losses corresponding to the loss of equation, boundary, and initial conditions, respectively. In addition, $L_{Eq}/L_{BC}/L_{IC}$ is computed following
$L_{Eq}$ $=\frac{1}{N}\sum^{N}_{i=1}|R_{i}(t,\bm{x},\bm{s})|^{2},$ (7)
$L_{BC}$ $=\frac{1}{N_{BC}}\sum^{N_{BC}}_{j=1}|e_{NN,j}^{eq}(t,\bm{x}_{w},\bm{s})-e_{j}(t,\bm{x}_{w},\bm{s})|^{2},$
$L_{IC}$ $=\frac{1}{N_{IC}}\sum^{N_{IC}}_{k=1}|e_{k}^{eq}(t_{0},\bm{x},\bm{s})-e_{NN,k}(t_{0},\bm{x},\bm{s})|^{2},$
@@
$R={\partial_{t}e^{\prime\prime}}_{NN}+\bm{s}\cdot\nabla e^{\prime\prime}_{NN}-\frac{1}{\rm{Kn}}(e^{eq}_{NN}-e^{\prime\prime}_{NN}),$ (8)
where $R$ denotes the residual of the equation, and $N$, $N_{BC}$ and $N_{IC}$ are the numbers of training points used to compute the loss functions in PINNs for the residual, boundary, and initial conditions, respectively; $R_{i}$, $e^{eq}_{j}/e_{NN,j}$ and $e^{eq}_{k}/e_{NN,k}$ are quantities evaluated at points $(t,\bm{x},\bm{s})_{i},i=1,...,N$, $(t,\bm{x}_{w},\bm{s})_{j},j=1,...,N_{BC}$ and $(t_{0},\bm{x},\bm{s})_{k},k=1,...,N_{IC}$, respectively. In the present study, we refer to the points used to evaluate $L_{Eq}$, $L_{B}C$, and $L_{IC}$ as residual, boundary, and initial points, respectively.
We note that the computations of the residual for the equation, i.e., $R_{i}$, require the calculation of the equilibrium energy density function, which is an integral defined in Eq. (2). Different approaches can be utilized to evaluate the integral since we can obtain the prediction of the energy density function at an arbitrary point $(t,\bm{x},\bm{s})$ in the temporal-spatial-angular space via the DNNs, e.g., Monte Carlo method, Gauss-Legendre, and Newton-Cotes quadrature. In the present study, it is found that Monte Carlo, which is computationally efficient at both low and high dimensions, is able to provide results with satisfactory accuracy. Hence, the Monte Carlo method is employed for the computations of $e^{eq}_{NN}$ in this work.
As for the selection of the training points for evaluating the loss function in PINNs (Eq. (7)), two different ways are often used, i.e., (1) the number ($N/N_{BC}/N_{IC}$) as well as the locations of the training points are prescribed at each training step, and (2) the number of points ($N/N_{BC}/N_{IC}$) is prescribed, but the corresponding locations are randomly generated at each training step. The former is similar to the meshes in the conventional numerical methods. As mentioned, lots of discrete points in the solid angular space are required to accurately capture the non-Fourier behavior at the ballistic regime, which results in expensive computational and memory costs for multiscale heat conduction problems in real-world applications. For the latter approach, if we randomly generate training points in the temporal-spatial-angular domain, denoted as $(t,\bm{x},\bm{s})_{i},i=1,...,M$, it is likely that each point $(t,\bm{x},\bm{s})_{i}$ will be unique, meaning that for different $i$, the points $(t,\bm{x})_{i}$ correspond to different angular direction $\bm{s})_{i}$. As a result, during the training process, we only update the energy density function at each $(t,\bm{x})_{i}$ based on one single angular direction $\bm{s}_{i}$. It can lead to slow convergence or even degenerated results since we need to compute the integral in the entire angular domain, as we will show in Sec. 3.1.

### v4

@@ Abstract
The phonon Boltzmann transport equation (BTE) is widely used for the description of multiscale heat conduction (from $nm$ to $\mu m$ or $mm$) in solid materials. Developing numerical approaches to solve this equation is challenging since it is a 7-dimensional integral-differential equation. In this work, we propose the Monte Carlo physics-informed neural networks (MC-PINNs), which do not suffer from the “curse of dimensionality”, to solve the phonon Boltzmann transport equation for modeling the multiscale heat conduction in solid materials. In MC-PINNs, we utilize a deep neural network to approximate the solution to the BTE and encode the BTE as well as the corresponding boundary/initial conditions using the automatic differentiation. In addition, we propose a novel two-step sampling approach to address the issues of inefficiency and inaccuracy in the widely used sampling methods in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial space (Step I) and then draw another number of points randomly in the solid angular space (Step II). The training points at each step are constructed based on the data drawn from the above two steps using the tensor product. The two-step sampling strategy enables the MC-PINNs (1) to model the heat conduction from ballistic to diffusive regimes, and (2) to be more memory efficient compared to the conventional numerical solvers or existing PINNs for BTE. A series of numerical examples including quasi-one-dimensional (quasi-1D) steady/unsteady heat conduction in a film, and the heat conduction in quasi-two- and three-dimensional square domains, are conducted to justify the effectiveness of the MC-PINNs for heat conduction spanning diffusive and ballistic regimes. Finally, we perform a comparison on the computational time and the memory usage between the MC-PINNs and one of the state-of-the-art numerical methods to demonstrate the potential of MC-PINNs for large-scale problems in real-world applications.
@@ 1 Introduction
Microdevices, e.g., microprocessors and microelectronic circuits, often generate heat with great density during operation. Efficient thermal management is crucial to the reliability and optimal performance of these devices. As reported in [1], the heat conduction in microdevices like chips is inherently a multiscale problem spanning nanoscale (1-100 nm) to microscale (0.1-100 $\mu$m) or macroscopic (0.1-100 mm). The classical Fourier’s law, which is a widely used empirical macroscopic model for heat conduction in solids, is not capable of describing the heat conduction in the non-diffusive or ballistic regime [2, 3, 4]. Nevertheless, the phonon Boltzmann transport equation (BTE) is a well-established model for the descriptions of the heat conduction in solid materials spanning the ballistic and diffusive regimes [5, 1, 6, 7].
@@
To advance the understanding of multiscale heat conduction in solid materials, lots of efforts have been taken into developing numerical methods for solving the phonon BTE. However, developing effective numerical methods to efficiently solve the phonon BTE is challenging since it is a 7-dimensional integral differential equation in real-world applications. The conventional numerical methods can be broadly categorized into two main approaches, i.e., the stochastic methods, e.g., direct simulation Monte Carlo (DSMC) [8, 9, 10, 11], etc., and deterministic approaches, e.g., discrete velocity/ordinate method [12, 13, 14, 15], synthetic iterative scheme [16, 17, 18], discrete unified gas kinetic scheme (DUGKS) [19, 20], and so on. The former uses discrete meshes in the temporal-spatial space and is efficient for steady problems in the ballistic regime. However, it converges slowly for problems in the (near) diffusive regime as well as non-stationary problems. The latter employs the discrete meshes in the temporal-spatial-angular space and is able to achieve good accuracy for steady/unsteady problems spanning a wide range of transport regimes. In general, the number of discrete points in the solid angular space required to obtain accurate results for the heat conduction problem in the diffusive regime is much smaller than in the ballistic regime. The requirement of large numbers of discrete points in the solid angular space to avoid the ray effect in the ballistic regime  [21, 22] leads to intensive computational cost and prohibitive memory usage for large-scale thermal engineering problems in practice.
@@
Deep neural networks (DNNs), which are able to break the curse of dimensionality, have recently emerged as an effective tool for solving partial differential equations (PDEs), such as physics-informed neural networks (PINNs) [23, 24, 25, 26], DeepRitz [27], and mothods combining moment closure with neural networks [28, 29, 30]. Generally, the DNN-based PDE solvers are mesh-free, which saves significant effort for generating the meshes in conventional methods, e.g., finite volume method, and finite element method. Among the aforementioned deep learning approaches, the PINNs, which encode the PDEs via the automatic differentiation technique, are one of the most widely used methods for solving PDE problems due to their effectiveness and straightforward implementations. Recently, the PINNs have been successfully employed to simulate flows ranging from continuous to rarefied flows by solving the Boltzmann-BGK or Boltzmann equation [31, 32], demonstrating their capability of handling multiscale problems [33, 34]. Inspired by the aforementioned work for modeling multiscale flows using PINNs, Li et al. proposed to solve the phonon BTE using PINNs to study the heat conduction in solid materials [35, 36]. The results showed that the PINNs are capable of modeling heat conduction ranging from ballistic to diffusive regimes with satisfactory accuracy. In particular, the solid angular space are discretized as a priori for training the PINNs in previous works [33, 35, 37]. For instance, the angular space is discretized utilizing the points in the Gauss Legendre quadrature rule in [35, 37], which is similar to the way that is widely used in the conventional deterministic approaches for solving phonon BTE, e.g., DUGKS. Although the effort of generating the meshes in temporal-spatial space can be saved in the PINN approaches, it shares the same defect as in the conventional deterministic approaches, i.e., a large number of discrete points in the angular space is required for heat conduction at the ballistic regime. Moreover, for scenarios in real-world applications where the heat conduction at the micro- and nanoscale coexists, a large number of discrete points in the angular space of BTE is also required for the accurate descriptions of the non-Fourier’s behavior at the ballistic regime, since it is challenging to develop adaptive discretization for heat conduction at different scales. Therefore, we still have difficulty employing the PINN methoeds developed in [35, 37] for large-scale heat conduction problems due to the expensive computational as well as memory cost. Numerical approaches for solving the phonon BTE that are more flexible as well as efficient are herein still desirable.
@@
In this study, we employ the PINNs as the backbone to solve the phonon BTE due to their effectiveness and capability for handling high-dimensional problems. In particular, we propose a novel two-step sampling approach for the training of PINNs to address the issue of expensive computational and memory costs in the existing PINN models. We refer to the PINNs equipped with the proposed two-step sampling approach as the Monte Carlo PINNs (MC-PINNs), which (1) is capable of modeling the heat conduction spanning ballistic to diffusive regimes, (2) is mesh-free in the temporal-spatial as well as the angular space, (3) is a unified approach at all regimes without the requirement to employ different discretizations in the angular space for heat conduction at different scales, and (4) requires much less discrete points in the angular space at each training step, and thus is more computationally and memory efficient comparing to the approach in [35, 37].
@@ 2.1 Phonon Boltzmann transport equation
The phonon BTE is widely used for simulating heat conduction in solid materials at different scales. Various models of the phonon BTE have been developed to address different applications [9]. For instance, in [38], the full/linearized phonon BTE is utilized to consider the complex interactions among different phonon modes; the frequency-resolved BTE is often employed in applications that require to consider waves with specific frequencies [10, 11, 17, 18]. Further, the gray model is generally used for multiscale heat conduction in three dimensions subject to minor temperature variations [20, 16, 18].
In this study, our particular interest is on the phonon gray model, which is specifically developed for probing multiscale heat conduction in three-dimensional materials with minor temperature variations. Given a certain number of reference variables, e.g., temperature $T_{ref}$, heat $C_{ref}$, length $L_{ref}$, group velocity $v_{g}$ and time $t_{ref}=L_{ref}/v_{g}$, the dimensionless phonon BTE in the gray model can then be obtained as
@@
where $e^{\prime\prime}$ represents the phonon distribution function of energy density, which is a function of time $t$, spatial coordinate $\bm{x}$, and the unit directional vector $\bm{s}$. The directional vector is uniformly distributed on the surface of a unit sphere in the phonon BTE, and thus can be expressed in the Cartesian coordinates as:
$\bm{s}=(s_{x},s_{y},s_{z})=(\cos\theta,\sin\theta\cos\varphi,\sin\theta\sin\varphi),$ (2)
where $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi]$ are the polar and azimuthal angles, respectively. The schematic of the directional vector in solid angular space with polar and azimuthal angle is shown in Fig. 1.
Figure 1: Schematic of the unit directional vector in solid angular space with polar angle denoting by the polar angle $\theta$ and azimuthal angle $\varphi$. $s_{x}$, $s_{y}$ and $s_{z}$ are the projections of the directional vector in Cartesian coordinates.
In addition, $e^{eq}$ denotes the equilibrium distribution function for the energy density; $C_{V}$ and $T$ are the specific heat and temperature, respectively. Moreover, $\text{Kn}=\lambda/L_{ref}$ is the Knudsen number defined as the ratio between the phonon mean free path to the characteristic length of a specific system. Due to the law of energy conservation, we can obtain the following equation for the right-hand side of Eq. (1a):
@@
where $d\Omega$ represents the integral over the surface of the unit sphere. The quantities of interest (QoI) in heat conduction problems including temperature $T$ and heat flux $\bm{q}$ can be obtained by taking the moment of phonon energy density distribution function, i.e.,
@@
Different boundary conditions are generally required to solve Eq. (1) in different applications. For example, the diffuse reflection conditions can be used for multiscale heat conduction with rough surfaces [16]; and the periodic boundary condition is often employed in scenarios with periodic geometries. In the current study, we adopt the thermalization boundary condition, in which we assume that the temperatures at the boundaries are known [39]. For this particular case, the phonons are absorbed upon striking the boundary, and new phonons with equilibrium states are emitted. We note that the aformentioned boundary condition is equivalent to imposing the Dirichlet boundary condition for the phonon energy density function $e^{\prime\prime}$, which is expressed as follows:
$e^{\prime\prime}(t,x_{w},\bm{s})=\frac{C_{V}T_{w}}{4\pi},\quad\bm{s}\cdot\bm{n}>0,$ (5)
where $x_{w}$ refers to the boundary, $\bm{n}$ is the unit normal vector pointing from the boundary to the computational domain, and $T_{w}$ denotes the temperature at the boundary. For initial conditions, we assume that the distribution function is at the equilibrium state with the temperature at the initial time, which reads as [20]:
$e^{\prime\prime}(t_{0},\bm{x},\bm{s})=\frac{C_{V}T_{0}}{4\pi},$ (6)
where $T_{0}$ denotes the temperature at the initial time $t_{0}$.
We would like to discuss that the multiscale heat conduction is generally divided into different transport regimes based on the Knudsen number, as shown in Fig. 2. For $\rm{Kn}<0.01$, the characteristic length of the system is much larger than the phonon mean free path, and the phonon transport is in the diffusive regime. The temperature field is generally described by the Fourier’s law. As $\rm{Kn}$ increases, the Fourier’s law begins to fail. For $\rm{Kn}>10$, the heat conduction is dominated by the ballistic effects.
Figure 2: Schematic of transport regimes based on the Knudsen number. $L$ is the characteristic length of the system and $\lambda$ is the phonon mean free path.
@@ 2.2.1 Physics-informed neural networks
The PINNs for solving phonon BTE are illustrated in Fig. 3. As shown, we first utilize DNNs, which take inputs as $(t,\bm{x},\bm{s})$, to approximate the solution to Eq. (1), i.e., $e^{\prime\prime}$, and then the governing equation, as well as the boundary and/or the initial conditions, are encoded in the DNNs via the automatic differentiation [40]. The loss function for training the PINNs is composed of the mean squared errors for the residual of the equation, mismatches between the predictions of DNNs and the boundary/initial conditions, which is expressed as follows:
$\mathcal{L}=\mathcal{L}_{Eq}+\mathcal{L}_{BC}+\mathcal{L}_{IC},$ (7)
where $\mathcal{L}$ is the total loss, $\mathcal{L}_{Eq}$, $\mathcal{L}_{BC}$, and $\mathcal{L}_{IC}$ denote the losses corresponding to the loss of equation, boundary, and initial conditions, respectively. Specifically, $\mathcal{L}_{Eq}$, $\mathcal{L}_{BC}$ and $\mathcal{L}_{IC}$ are computed following
$\mathcal{L}_{Eq}$ $=\frac{1}{N}\sum^{N}_{i=1}|R_{i}(t,\bm{x},\bm{s})|^{2},$ (8)
$\mathcal{L}_{BC}$ $=\frac{1}{N_{BC}}\sum^{N_{BC}}_{j=1}|e_{NN,j}(t,\bm{x}_{w},\bm{s})-e^{eq}_{j}(t,\bm{x}_{w},\bm{s})|^{2},$
$\mathcal{L}_{IC}$ $=\frac{1}{N_{IC}}\sum^{N_{IC}}_{k=1}|e_{NN,k}(t_{0},\bm{x},\bm{s})-e_{k}^{eq}(t_{0},\bm{x},\bm{s})|^{2},$
@@
$R={\partial_{t}e}_{NN}+\bm{s}\cdot\nabla e_{NN}-\frac{1}{\rm{Kn}}(e^{eq}-e_{NN}),$ (9)
where $R$ denotes the residual of the equation, $N$, $N_{BC}$ and $N_{IC}$ are the numbers of training points used to compute the loss functions in PINNs for the residual, boundary, and initial conditions, respectively; $R_{i}$, $e^{eq}_{j}/e_{NN,j}$ and $e^{eq}_{k}/e_{NN,k}$ are quantities evaluated at points $(t,\bm{x},\bm{s})_{i}$ for $i=1,...,N$; $(t,\bm{x}_{w},\bm{s})_{j}$ for $j=1,...,N_{BC}$; and $(t_{0},\bm{x},\bm{s})_{k}$ for $k=1,...,N_{IC}$, respectively. In the present study, we refer to the points used to evaluate $\mathcal{L}_{Eq}$, $\mathcal{L}_{BC}$, and $\mathcal{L}_{IC}$ as residual, boundary, and initial points, respectively.
We note that the computations of the residual for the equation, i.e., $R_{i}$, require the calculation of the equilibrium energy density function, which is an integral defined in Eq. (3). Different approaches can be utilized to compute the integral since we can obtain the prediction of the energy density function at an arbitrary point $(t,\bm{x},\bm{s})$ in the temporal-spatial-angular space via the DNNs, e.g., Monte Carlo method, Gauss-Legendre, and Newton-Cotes quadrature. In the present study, it is found that Monte Carlo method, which is computationally efficient at both low and high dimensions, is able to provide results with satisfactory accuracy. Hence, the Monte Carlo method is employed for the computations of $e^{eq}$ in this work.
As for the selection of the training points for evaluating the loss function in PINNs (Eq. (8)), two different ways are often used, i.e., (1) the number ($N/N_{BC}/N_{IC}$) as well as the locations of the training points are prescribed at each training step, and (2) the number of points ($N/N_{BC}/N_{IC}$) is prescribed, but the corresponding locations are randomly generated at each training step. The former is similar to the meshes in the conventional numerical methods. As mentioned, lots of discrete points in the solid angular space are required to accurately capture the non-Fourier behavior at the ballistic regime, which results in expensive computational and memory costs for multiscale heat conduction problems in real-world applications. For the latter approach, if we randomly generate training points in the temporal-spatial-angular domain, denoted as $(t,\bm{x},\bm{s})_{i},i=1,...,M$, it is probable that each point $(t,\bm{x},\bm{s})_{i}$ will be different for different $i$. As a result, we only solve the energy density function corresponds to $\bm{s}_{i}$ at $(t,\bm{x})_{i}$ in each training step. It leads to slow convergence or even degenerated results since we need to compute the integral in the entire angular domain, as we will show in Sec. 3.1.